# Provable Robustness in Multimodal Large Language Models via Feature Space Smoothing

### v1

@@ Abstract
Multimodal large language models (MLLMs) exhibit strong capabilities across diverse applications, yet remain vulnerable to adversarial perturbations that distort their feature representations and induce erroneous predictions. To address this vulnerability, we propose the Feature-space Smoothing (FS) and theoretically prove that FS offers certified robustness on the feature representations of MLLMs. Specifically, FS transforms any feature encoder into a smoothed variant that is guaranteed to maintain a certified lower bound on the feature cosine similarity between clean and adversarial representations under $\ell_{2}$-bounded attacks. Moreover, we indicate that the value of this Feature Cosine Similarity Bound (FCSB) derived from FS can be improved by enlarging the defined Gaussian robustness score on the vanilla encoder. Building upon this, we introduce the Purifier and Smoothness Mapper (PSM), a plug-and-play module that improves the Gaussian robustness score of MLLMs and thus enhances their certified robustness under FS, without requiring any retraining on MLLMs. We demonstrate that the FS with PSM not only provides a strong theoretical robustness guarantee but also exhibits superior empirical performance compared to adversarial training. Extensive experiments across diverse MLLMs and downstream tasks indicate the effectiveness of the FS-PSM, reducing the Attack Success Rate (ASR) of various white-box attacks from nearly 90% to about 1%.
@@ 1 Introduction
The emergence of the Multimodal Large Language Models (MLLMs), such as GPT-5, Gemini 2.5, and Claude 3.7 Sonnet, has fundamentally reshaped existing working paradigms and significantly advanced societal productivity. Despite their remarkable capabilities across a broad spectrum of real-world tasks, these models still encounter critical safety challenges, such as adversarial vulnerabilities [51, 5, 23, 14].
@@
Figure 1: Illustration of the FS-PSM, which guarantees that the cosine similarity of the adversarial and clean features extracted by MLLM’s encoder is larger than FCSB for robust predictions.
@@
Adversaries can manipulate predictions of the MLLMs to a malicious state by injecting subtle and imperceptible perturbations to the clean inputs, exploiting the models’ insufficient local smoothness and uncontrolled Lipschitz continuity [10, 4, 12, 47].
@@
Countermeasures towards those threats can be roughly classified into empirical defense and certified defense. Typical empirical approaches include adversarial training [27, 35, 41, 45, 2, 37, 28, 29] and input purification [27, 35, 41, 45, 2, 37, 28, 29]. Despite their demonstrated empirical effectiveness, the multimodal nature of MLLMs also poses a great challenge for existing adversarial defense methods. Since these models accept heterogeneous inputs across diverse domains, ensure a robust encoder via adversarial training that can generalize to various scenarios is challenging and computationally expensive. Moreover, these approaches lack formal robustness guarantees and remain susceptible to stronger or adaptive adversaries [38, 3, 44]. In contrast, certified approaches aim to guarantee that the model returns a constant prediction result within a certain range, usually a $\ell_{2}$ or $\ell_{\infty}$-norm constrained area [34, 43, 11, 15, 4, 47, 36]. However, most previous certified defense approaches [4, 47, 42, 11, 36] predominantly assume the prediction of one-dimensional outputs (e.g., class label), thereby limiting their applicability to more general tasks on MLLMs.
@@
To address the aforementioned challenges, we propose the Feature-space Smoothing (FS), a provable defense method that offers certified robustness for the feature representations of MLLMs. FS smooths the vanilla feature encoder of the MLLM, and the resulted smoothed encoder is guaranteed to maintain a provable lower bound on the cosine similarity between clean and adversarial representations under $\ell_{2}$-norm bounded attacks.
@@
Moreover, we demonstrate that the Feature Cosine Similarity Bound (FCSB) of the smoothed feature encoder is intrinsically determined by a Gaussian robustness score, which is defined to measure the prediction consistency of the vanilla feature extractor under Gaussian noise. Nonetheless, without Gaussian noise–augmented training, the Gaussian robustness of these MLLMs remains limited, leading to a correspondingly suboptimal FCSB. To address this and avoid the heavy computational cost of fine-tuning MLLMs, we propose a plug-and-play Purifier and Smoothness Mapper (PSM) that effectively enhance the Gaussian robustness score for the vanilla feature encoder in a training-free way. Specifically, the purifier operates prior to feature extraction and is trained to denoise Gaussian perturbations and maximize the Gaussian robustness score. Meanwhile, the smoothness mapper performs post-extraction refinement and is trained to preserve the feature distribution while further enhancing this score without fine-tuning the encoder. These two components of PSM work synergistically to enhance the certified robustness of MLLMs under FS.
@@
To implement our PSM, we employ a pre-trained guided diffusion model as the Gaussian noise purifier and design a noise-aware residual network as the smoothness mapper. The PSM is optimized using the proposed utility–robustness loss and trained on data from diverse visual domains to enhance the Gaussian robustness while preserving the feature utility for the encoder. To comprehensively assess the performance, we evaluate our method against state-of-the-art (SOTA) adversarial attacks tailored for MLLMs under the white-box setting and compare it with strong adversarial training defense. Experimental results demonstrate that our FS with PSM enhanced provides strong protection for various MLLMs under diverse downstream tasks. Overall, the main contributions of this work are:
@@
  • We propose the Feature-space Smoothing (FS) to turn any feature encoder of MLLMs into a smoothed version, and theoretically prove that the smoothed encoder maintains a certified lower bound on the feature cosine similarity between clean and adversarial representations.
@@
  • We propose a plug-and-play Purifier and Smoothness Mapper (PSM) that effectively enhances the value of the FCSB of the smoothed encoder in a training-free manner.
@@
  • We conduct extensive experiments demonstrating that integrating our proposed FS-PSM greatly enhances the adversarial performance of various MLLMs and dramatically reduces the ASR under various white-box attacks.
@@ 2 Related Work
Adversarial attacks on MLLMs. While MLLMs continue to achieve remarkable performance across diverse applications, extensive works [31, 5, 51, 14, 23, 39, 49, 50, 48] have exposed their adversarial vulnerabilities, raising serious safety concerns. Early work, such as AttackVLM [51], explores transferable attacks by disrupting the feature representations of CLIP [33] and BLIP [19], showing strong adversarial transferability among open-source models but limited effectiveness against closed-source commercial systems. More recent approaches, such as M-Attack [23] and FOA-Attack [14], further advance this direction by leveraging multi-extractor ensembles and feature-space alignment, achieving over 90% targeted attack success rates on image-captioning tasks against powerful closed-sourced commercial MLLMs (e.g., ChatGPT-4o). This highlights an urgent need for trustworthy defenses that provide effective and provable protection for MLLMs.
Adversarial defense on MLLMs. Adversarial defense methods can be broadly classified into empirical and provable approaches. Empirical defenses for MLLMs mainly include adversarial training [27, 35, 41, 45, 2, 37, 28, 29], which enhances robustness by augmenting training data with adversarial examples, and input purification [30, 17, 22], which employs generative mechanisms such as diffusion models or autoencoders to recover clean inputs prior to inference. Recent studies [28, 29, 37] have revealed that utilization of adversarially trained CLIP-feature encoders can enhance adversarial robustness for MLLMs. However, these methods lack formal robustness guarantees and remain vulnerable to adaptive and unseen threats [38, 3, 44]. Moreover, adversarial training demands costly retraining and often leads to degradation in clean performance.
Alternatively, certified defenses aim to provide mathematically provable robustness guarantees. The use of Gaussian smoothing for certified robustness was initially introduced for classification models [4, 18, 16], yet its theoretical formulation is restricted to one-dimensional outputs, limiting its applicability to tasks such as auto-regression or multimodal generation. To overcome these limitations and ensure trustworthy protection for MLLMs, we propose the Feature-space Smoothing (FS), a defense mechanism that establishes provable adversarial robustness for MLLMs.
3 Feature Space Smoothing
@@ 3.1 Preliminary
$\vskip-2.84526pt\mathop{\max}\limits_{{{\bm{x}}}+{\bm{\delta}}\in\mathcal{B}_{\epsilon}({\bm{x}})}\mathcal{L}\left({\mathcal{F}\left({{{\bm{x}}}}\right),\mathcal{F}\left({{{\bm{x}}}+{\bm{\delta}}}\right)}\right).\vskip-2.84526pt$ (1)
Adversarial effects on feature representations: While adversarial attacks primarily aim to alter the model’s predictions, numerous studies [14, 46, 20, 13, 21, 23, 40, 32, 8] have shown that successful attacks typically induce substantial distortions in the model’s feature representations. Let $\bm{x}$$^{\prime}$ denote the adversarial example with adversarial feature $\bm{z}$$^{\prime}$. Let ${\bm{z}}^{t}$ be the adversarial targeted feature with malicious semantic meaning. The attack generally leads to $\max\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}}\right)$ for untargeted attacks and $\min\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}^{t}}\right)$ for targeted attacks. Thus, ensuring a robust feature encoder that $\min\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}}\right)$ is crucial for the trustworthy prediction.
Randomized smoothing: Consider a $k$ classes classification problem with the input ${\bm{x}}\in{\mathbb{R}^{d}}$ and the label ${\bm{y}}\in\mathcal{Y}=\left\{{{c_{1}},\ldots,{c_{k}}}\right\}$. Randomized Smoothing (RS) first corrupts each input $\bm{x}$ by adding the Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(0,{\sigma^{2}}\bm{I})$. Then it turns an arbitrary base classifier $\mathcal{F}$ into a smoothed version $\hat{\mathcal{F}}$ that possesses $\ell_{2}$} certified robustness guarantees. The smoothed classifier $\hat{\mathcal{F}}$ returns whichever the class the base classifier $\mathcal{F}$ is most likely to return among the distribution ${\bm{x}}+\bm{\varepsilon}\sim\mathcal{N}({\bm{x}},{\sigma^{2}}\bm{I})$, which is:
$\begin{array}[]{l}\hat{\mathcal{F}}({\bm{x}})=\mathop{\arg\max}\limits_{c\in\mathcal{Y}}{\mathbb{P}}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c).\vskip-5.69054pt\end{array}$ (2)
@@ 3.2 Certified Bound via Feature Space Smoothing
where $I$ denotes the $d\times d$ identity and $\|\cdot\|$ is the Euclidean norm. $\bm{\varepsilon}\sim\mathcal{N}(0,I)$ denotes Gaussian noise with zero mean and standard deviation one. The smoothed feature encoder outputs the expectation of feature representations over the Gaussian distribution $\mathcal{N}({\bm{x}},I)$.
Figure 2: The training framework of the PSM. The purifier performs pre-processing and the smoothness mapper refines post-extracted features to enhance the Gaussian robustness. Parameters of MLLMs are frozen, and the purifier and mapper are optimized with $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\mathcal{M}}$. For evaluation, the input $\bm{x}$} is replaced with adversarial input $\bm{x}$$^{\prime}$ and the forward calculation marked by orange color will be removed.
@@
$\vskip-5.69054ptS_{{\bm{x}}_{t}}({\bm{x}})=\tfrac{1}{2}\Big(1+\mathrm{Cos}\big(f_{e}({\bm{x}}),\,f_{e}({\bm{x}}_{t})\big)\Big)$ (4)
@@
The score $\hat{S}({\bm{x}})$ evaluates the expected cosine similarity between the feature representation of a Gaussian-perturbed input ${\bm{x}}+\bm{\varepsilon}$ and that of the clean input $\bm{x}$}, characterizing the Gaussian robustness of the vanilla feature encoder ${f}_{e}({\bm{x}})$. Notably, we prove that this score $\hat{S}({\bm{x}})$ preserves a good Lipschitz property, which serves as a theoretical foundation for proving the certified robustness for $\hat{f}_{e}({\bm{x}})$.
Lemma 1 (Lipschitz property for the Gaussian robustness score).
@@
The proof of Lemma 1 is in the supplementary material, Section S.1. It implies that the mapping from $\bm{x}$} to $\Phi^{-1}(\hat{S}({\bm{x}}))$ exhibits strong adversarial robustness, as it satisfies a $1$-Lipschitz constraint. Generally, $\hat{S}({\bm{x}})$ measures the Gaussian robustness of the vanilla feature encoder $f_{e}(x)$. We then prove that this score $\hat{S}({\bm{x}})$ fundamentally determines the value of the certified robustness bound of its smoothed encoder $\hat{f}_{e}(x)$. The following Theorem establishes an explicit relationship between this score and the lower bound on the adversarial feature cosine similarity.
Theorem 1 (Certified lower bound on the adversarial feature cosine similarity).
@@
Denote $2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$ as the Feature Cosine Similarity Bound (FCSB). Theorem 1 reveals that:
@@
The proof of Theorem 1 is in the supplementary material, Section S.1.
Corollary 1 (Certified radius $\mathcal{R}$ for adversarial cosine similarity $\geq$ 0.5).
@@
Building upon Theorem 1, Corollary 1 establishes a certified radius $\mathcal{R}$ for the smoothed feature encoder. For any adversarial perturbation satisfying $\|\delta\|_{2}\leq\mathcal{R}$, the FCSB of the smoothed encoder is guaranteed to remain above 0.5.
Why feature-space smoothing for MLLMs? Unlike previous smoothing methods, performing smoothing in the feature space offers several advantages:
  • Efficiency: Compared to the previous RS that smooths the entire model $\mathcal{F}$, smoothing the feature encoder $f_{e}$, which is substantially lighter than the whole MLLM $\mathcal{F}$, significantly reduces the computational time of multiple forward calculations required by smoothing.
  • Generality: The FS provides certified robustness at the feature representation level, making it applicable to various downstream tasks (e.g. image captioning, classification and visual question answering).
  • Effectiveness: Since feature representations play a critical role in the final prediction, guaranteeing a trustworthy feature representation effectively improves the prediction reliability and robustness under adversarial attacks.
4 Purifier and Smoothness Mapper
Theorem 1 and Corollary 1 reveal an intriguing robustness property of the smoothed feature encoder $\hat{f}_{e}$. However, MLLMs’ feature encoders $f_{e}$ generally exhibits limited Gaussian robustness, which restricts the value of FCSB derived on $\hat{f}_{e}$. One solution is to estimate the smoothness score $\hat{S}({\bm{x}})$ through Monte Carlo sampling and training $f_{e}$ to maximize this score via gradient backpropagation. However, this could reduce the adaptability and practicality of our FS protection, as fine-tuning and re-aligning the LLM with multi-modal feature encoders is highly complex and costly. This thereby motivates us to propose the Purifier and Smoothness Mapper (PSM), a plug-and-play module that can be seamlessly integrated with MLLMs to enhance its Gaussian robustness score $\hat{S}(x)$.
The training framework of our proposed PSM is shown in Figure 2. The purifier $\mathcal{P}$ operates before feature extraction to denoise the Gaussian perturbations, and the smoothness mapper $\mathcal{M}$ performs post-extraction to do the feature refinement. Those two modules work together to enhance the Gaussian robustness of the given feature encoder.
4.1 Noise purifier
To denoise the Gaussian noise, $\mathcal{P}$ is trained to minimize the reconstruction loss $l_{mse}$, defined as:

### v2

# Feature-Space Adversarial Robustness Certification for Multimodal Large Language Models

@@ Abstract
Multimodal large language models (MLLMs) exhibit strong capabilities across diverse applications, yet remain vulnerable to adversarial perturbations that distort their feature representations and induce erroneous predictions. To address this vulnerability, we propose Feature-space Smoothing (FS), a general framework that provides certified robustness guarantees at the feature representation level of MLLMs. We theoretically prove that FS converts a given feature extractor into a smoothed variant that is guaranteed a certified lower bound on the cosine similarity between clean and adversarial features under $\ell_{2}$-bounded perturbations. Moreover, we establish that the value of this Feature Cosine Similarity Bound (FCSB) is determined by the intrinsic Gaussian robustness score of the given encoder. Building on this insight, we introduce the Gaussian Smoothness Booster (GSB), a plug-and-play module that enhances the Gaussian robustness score of pretrained MLLMs, thereby strengthening the robustness guaranteed by FS, without requiring additional MLLM retraining. Extensive experiments demonstrate that applying the FS to various MLLMs yields strong certified feature-space robustness and consistently leads to robust task-oriented performance across diverse applications.
Machine Learning, ICML
@@ 1 Introduction
The emergence of the Multimodal Large Language Models (MLLMs), such as GPT-5 (OpenAI, 2025), Gemini 3 Pro (DeepMind, 2025), and Claude Sonnet 4.5 (Anthropic, 2025), has fundamentally reshaped existing working paradigms and significantly advanced societal productivity. Despite their remarkable capabilities across a broad spectrum of real-world tasks, these models still encounter critical safety challenges, such as adversarial vulnerabilities (Zhao et al., 2023; Cui et al., 2024; Li et al., 2025b; Jia et al., 2025).
@@
Figure 1: Illustration of the FS-GSB, which guarantees that the cosine similarity of the adversarial and clean features extracted by MLLM’s encoder is larger than FCSB for robust predictions.
@@
Adversaries can manipulate predictions of the MLLMs to a malicious state by injecting subtle and imperceptible perturbations to the clean inputs, exploiting the models’ insufficient local smoothness and uncontrolled Lipschitz continuity (Goodfellow et al., 2014; Cohen et al., 2019; Hein and Andriushchenko, 2017; Xia et al., 2024b).
@@
Countermeasures towards those threats can be roughly classified into empirical defense and certified defense. Typical empirical approaches include adversarial training (Madry et al., 2018; Rebuffi et al., 2021; Wang et al., 2024c; Xhonneux et al., 2024; Casper et al., 2024; Schlarmann et al., 2024; Malik et al., 2025; Mao et al., 2023) and input purification (Nie et al., 2022; Yoon et al., 2021; Lee and Kim, 2023; Lei et al., 2025; Zollicoffer et al., 2025). Despite their demonstrated empirical effectiveness, these approaches lack formal robustness guarantees and remain susceptible to stronger adversaries (Tramer et al., 2020; Chen et al., 2023; Wu et al., 2020; Olivier and Raj, 2023). Moreover, the multimodal nature of MLLMs also poses a great challenge for existing adversarial training methods. Since they accept heterogeneous inputs across diverse domains, ensure a robust encoder via adversarial training that can generalize to various scenarios is challenging and computationally costly.
In contrast, certified approaches aim to guarantee that the model returns a constant prediction result within a certain range, usually a $\ell_{2}$ or $\ell_{\infty}$-norm constrained area (Raghunathan et al., 2018; Wong and Kolter, 2018; Hao et al., 2022; Kakizaki et al., 2023; Cohen et al., 2019; Xia et al., 2024b; Salman et al., 2019). However, most previous certified defense approaches (Cohen et al., 2019; Xia et al., 2024b; Wang et al., 2024b; Hao et al., 2022; Salman et al., 2019) predominantly assume the prediction of one-dimensional outputs (e.g., class label), thereby limiting their applicability to more general tasks on MLLMs.
@@
To address the aforementioned challenges, we propose the Feature-space Smoothing (FS), a provable defense method that offers certified robustness on the feature representations of MLLMs. Specifically, FS smooths the vanilla feature encoder of the MLLM, and the resulted smoothed encoder is guaranteed to maintain a provable lower bound on the cosine similarity between clean and adversarial representations under $\ell_{2}$-norm bounded attacks. Moreover, we demonstrate that this Feature Cosine Similarity Bound (FCSB) of the smoothed feature encoder is intrinsically determined by a defined Gaussian robustness score, which measures the prediction consistency of the vanilla feature extractor under Gaussian noise.
@@
Nonetheless, the Gaussian robustness of these MLLMs remains limited without Gaussian noise augmented training, leading to a correspondingly suboptimal FCSB. To address this whiling avoiding the high computational cost of fine-tuning MLLMs, we propose a plug-and-play Gaussian Smoothness Booster (GSB), consisting of a lightweight Gaussian denoiser and a residual smoothness mapper. Specifically, the denoiser is applied prior to the feature encoder and is trained to mitigates Gaussian noises and improves the resulting robustness score. In parallel, the smoothness mapper operates after feature extraction for feature-level refinement. It is trained to preserve the original feature distribution while further enhancing the Gaussian robustness score. These two components of GSB work synergistically to enhance the certified robustness of MLLMs provided by FS. The GSB is optimized using the proposed utility–robustness loss and trained on data from diverse visual domains to enhance the Gaussian robustness while preserving the feature utility for the encoder.
@@
To comprehensively assess the performance, we evaluate our method against state-of-the-art (SOTA) adversarial attacks tailored for MLLMs under the white-box setting and compare it with advanced adversarial training methods. Extensive experimental results demonstrate that our FS not only provides a strong certified robustness guarantee on the feature representation but also exhibits superior adversarial robustness for various MLLMs under diverse downstream tasks. Overall, the contributions of this work are:
@@
  • We propose the Feature-space Smoothing (FS) to turn any feature encoder into a smoothed version, and theoretically prove that the smoothed encoder maintains a certified lower bound on the feature cosine similarity between clean and adversarial representations.
@@
  • We propose the Gaussian Smoothness Booster (PSB) that effectively enhances the FCSB of the smoothed encoder in MLLMs via a plug-and-play manner.
@@
  • We conduct extensive experiments demonstrating that integrating our proposed FS-GSB greatly enhances the adversarial robustness of various MLLMs and dramatically reduces the ASR under various white-box attacks.
@@ 2 Related Work
Adversarial attacks on MLLMs. While MLLMs continue to achieve remarkable performance across diverse applications, extensive works (Qi et al., 2024; Cui et al., 2024; Zhao et al., 2023; Jia et al., 2025; Li et al., 2025b; Wang et al., 2024a; Zhang et al., 2024, 2025; Xie et al., 2025) have exposed their adversarial vulnerabilities, raising serious safety concerns. Early work, such as AttackVLM (Zhao et al., 2023), explores transferable attacks by disrupting the feature representations of CLIP (Radford et al., 2021) and BLIP (Li et al., 2023a), showing strong adversarial transferability among open-source models but limited effectiveness against closed-source commercial systems. More recent approaches, such as M-Attack (Li et al., 2025b) and FOA-Attack (Jia et al., 2025), further advance this direction by leveraging multi-extractor ensembles and feature-space alignment, achieving over 90% targeted attack success rates on image-captioning tasks against powerful closed-sourced commercial MLLMs (e.g., ChatGPT-4o). This highlights an urgent need for trustworthy defenses that provide effective and provable protection for MLLMs.
Adversarial defense on MLLMs. Adversarial defense methods can be broadly classified into empirical and provable approaches. Empirical defenses for MLLMs mainly include adversarial training (Madry et al., 2018; Rebuffi et al., 2021; Wang et al., 2024c; Xhonneux et al., 2024; Casper et al., 2024; Schlarmann et al., 2024; Malik et al., 2025; Mao et al., 2023), which enhances robustness by augmenting training data with adversarial examples, and input purification (Nie et al., 2022; Lei et al., 2025; Li et al., 2025a), which employs generative mechanisms such as diffusion models or autoencoders to recover clean inputs prior to inference. Recent studies (Malik et al., 2025; Mao et al., 2023; Schlarmann et al., 2024) have revealed that utilization of adversarially trained CLIP-feature encoders can enhance adversarial robustness for MLLMs. However, these methods lack formal robustness guarantees and remain vulnerable to adaptive and unseen threats (Tramer et al., 2020; Chen et al., 2023; Wu et al., 2020). Moreover, adversarial training demands costly retraining and often leads to degradation in clean performance.
Alternatively, certified defenses aim to provide mathematically provable robustness guarantees. The use of Gaussian smoothing for certified robustness was initially introduced for classification models (Cohen et al., 2019; Li et al., 2018; Lecuyer et al., 2019), yet its theoretical formulation is restricted to one-dimensional outputs, limiting its applicability to tasks such as auto-regression or multimodal generation. To overcome these limitations and ensure trustworthy protection for MLLMs, we propose the Feature-space Smoothing (FS), a defense mechanism that establishes provable adversarial feature robustness for MLLMs.
3 Feature-Space Smoothing
@@ 3.1 Preliminary
$\vskip-2.84526pt\mathop{\max}\limits_{{{\bm{x}}}+{\bm{\delta}}\in\mathcal{B}_{\epsilon}({\bm{x}})}\mathcal{L}\left({\mathcal{F}\left({{{\bm{x}}}}\right),\mathcal{F}\left({{{\bm{x}}}+{\bm{\delta}}}\right)}\right).$ (1)
Adversarial effects on feature representations: While adversarial attacks primarily aim to alter the model’s predictions, numerous studies (Jia et al., 2025; Xia et al., 2024a; Li et al., 2020; Huang et al., 2019; Li et al., 2023b; Ding et al., 2024) have shown that successful attacks typically induce substantial distortions in the model’s feature representations. Let $\bm{x}$$^{\prime}$ denote the adversarial example with adversarial feature $\bm{z}$$^{\prime}$. Let ${\bm{z}}^{t}$ be the adversarial targeted feature with malicious semantic meaning. The attack generally leads to $\max\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}}\right)$ for untargeted attacks and $\min\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}^{t}}\right)$ for targeted attacks. Thus, ensuring a robust feature encoder that $\min\mathcal{L}\left({{\bm{z}}^{\prime},{\bm{z}}}\right)$ is crucial for the trustworthy prediction.
Randomized smoothing: Consider a $k$ classes classification problem with the input ${\bm{x}}\in{\mathbb{R}^{d}}$ and the label ${\bm{y}}\in\mathcal{Y}=\left\{{{c_{1}},\ldots,{c_{k}}}\right\}$. Randomized Smoothing (RS) first corrupts each input $\bm{x}$ by adding the Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(0,{\sigma^{2}}\bm{I})$. Then it turns an arbitrary base classifier $\mathcal{F}$ into a smoothed version $\hat{\mathcal{F}}$ that possesses $\ell_{2}$} certified robustness guarantees. The smoothed classifier $\hat{\mathcal{F}}$ returns whichever class the base classifier $\mathcal{F}$ is most likely to return among the distribution ${\bm{x}}+\bm{\varepsilon}\sim\mathcal{N}({\bm{x}},{\sigma^{2}}\bm{I})$, which is:
$\begin{array}[]{l}\hat{\mathcal{F}}({\bm{x}})=\mathop{\arg\max}\limits_{c\in\mathcal{Y}}{\mathbb{P}}(\mathcal{F}({\bm{x}}+\bm{\varepsilon})=c).\vskip-2.84526pt\end{array}$ (2)
@@ 3.2 Certified Bound via Feature Space Smoothing
where $I$ denotes the $d\times d$ identity and $\|\cdot\|$ is the Euclidean norm. $\bm{\varepsilon}\sim\mathcal{N}(0,I)$ denotes Gaussian noise with zero mean and standard deviation one. The smoothed feature encoder $\hat{f}_{e}(\bm{x})$ outputs the expectation of feature representations over the Gaussian distribution $\mathcal{N}({\bm{x}},I)$.
Figure 2: The training framework of the GSB. The denoiser performs pre-processing, and the smoothness mapper refines post-extracted features to enhance the Gaussian robustness. Parameters of MLLMs are frozen, and the denoiser and mapper are optimized with $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\mathcal{M}}$. For evaluation, $\bm{x}$} is replaced with adversarial input $\bm{x}$$^{\prime}$, and the forward calculation marked in orange color will be removed.
@@
$\vskip-2.84526ptS_{{\bm{x}}_{t}}({\bm{x}})=\tfrac{1}{2}\Big(1+\mathrm{Cos}\big(f_{e}({\bm{x}}),\,f_{e}({\bm{x}}_{t})\big)\Big)$ (4)
@@
The score $\hat{S}({\bm{x}})$ evaluates the expected cosine similarity between the feature representation of a Gaussian-perturbed input ${\bm{x}}+\bm{\varepsilon}$ and that of the clean input $\bm{x}$}, characterizing the Gaussian robustness of the vanilla feature encoder ${f}_{e}$. Notably, we prove that this score $\hat{S}({\bm{x}})$ preserves a good Lipschitz property, which serves as a theoretical foundation for proving the certified robustness for $\hat{f}_{e}({\bm{x}})$.
Lemma 3.1 (Lipschitz property for the Gaussian robustness score).
@@
The proof of Lemma 3.1 is in the Appendix, Section A.1.1. It implies that the mapping from $\bm{x}$} to $\Phi^{-1}(\hat{S}({\bm{x}}))$ exhibits strong adversarial robustness, as it satisfies a $1$-Lipschitz constraint. Generally, $\hat{S}({\bm{x}})$ measures the Gaussian robustness of the vanilla feature encoder $f_{e}(x)$. We then prove that this score $\hat{S}({\bm{x}})$ fundamentally determines the value of the certified robustness bound of its smoothed encoder $\hat{f}_{e}(x)$. The following Theorem establishes an explicit relationship between this score and the lower bound on the adversarial feature cosine similarity.
Theorem 3.2 (Certified lower bound on the adversarial feature cosine similarity).
@@
Denote $2{\Phi}\left({\Phi^{-1}}(\hat{S}({\bm{x}}))-\epsilon\right)-1$ as the Feature Cosine Similarity Bound (FCSB). Theorem 3.2 reveals that:
@@
The proof of Theorem 3.2 is in the Appendix, Section A.1.2.
Corollary 3.3 (Certified radius $\mathcal{R}$ for adversarial cosine similarity $\geq$ 0.5).
@@
Building upon Theorem 3.2, Corollary 3.3 establishes a certified radius $\mathcal{R}$ for the smoothed feature encoder. For any adversarial perturbation satisfying $\|\delta\|_{2}\leq\mathcal{R}$, the FCSB of the smoothed encoder is guaranteed to remain above 0.5.
4 Gaussian Smoothness Booster
Theorem 3.2 and Corollary 3.3 reveal an intriguing robustness property of the smoothed feature encoder $\hat{f}_{e}$. However, feature encoders $f_{e}$ of MLLMs generally exhibit limited Gaussian robustness, which restricts the value of FCSB derived on $\hat{f}_{e}$. One solution is to estimate the smoothness score $\hat{S}({\bm{x}})$ through Monte Carlo sampling and training $f_{e}$ to maximize this score via gradient backpropagation. However, this could reduce the adaptability and practicality of the proposed FS framework, as fine-tuning and re-aligning the LLM is highly complex and costly. This thereby motivates us to propose the Gaussian Smoothness Booster (GSB), a plug-and-play module that can be seamlessly integrated with MLLMs to enhance its Gaussian robustness score $\hat{S}(x)$.
The training framework of our proposed GSB is shown in Figure 2. The denoiser $\mathcal{P}$ operates before feature extraction to denoise the Gaussian perturbations, and the smoothness mapper $\mathcal{M}$ performs post-extraction to do the feature refinement. Those two modules work together to enhance the Gaussian robustness of the given feature encoder.
4.1 Gaussian Denoiser
To perform the Gaussian denoising, the denoiser $\mathcal{P}$ is trained to minimize the reconstruction loss $l_{mse}$, defined as: